OpenStudy (anonymous):

What is the value of a in the equation 4a + 3(a + 1) = 24? (4 points) 2 21 7 3

OpenStudy (anonymous):

@Holly00d1248 I think it is 2 but im not sure

OpenStudy (anonymous):

Okay thanks!

OpenStudy (mathstudent55):

Solve the equation. First, use the distributive property on the left side.

OpenStudy (anonymous):

4a+3a+3=24 7a=24-3 a=21/7=3

OpenStudy (anonymous):

That makes sence

OpenStudy (anonymous):

can you help with 2 others?

OpenStudy (anonymous):

yes i can try

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

An equation is shown below: 3(4x + 5) = 12 Which of the following correctly shows the beginning steps to solve this equation? (4 points) Step 1: 7x + 5 = 12 Step 2: 7x = 7 Step 1: 12x + 5 = 12 Step 2: 12x = 7 Step 1: 12x + 15 = 12 Step 2: 12x = –3 Step 1: 7x + 8 = 12 Step 2: 7x = 4

OpenStudy (mathstudent55):

Do you know how to use the distributive property?

OpenStudy (mathstudent55):

For example: \(3(4x + 6) = 3 \cdot 4x + 3 \cdot 6 = 12x + 18\)

OpenStudy (anonymous):

How many solutions can be found for the equation 5r + 10 – 5r = 10? (4 points) Zero One Infinitely many Two

OpenStudy (mathstudent55):

You have 3(4x + 5) on the left side. Follow the example above, and use the distributive property on the left side. What do you get?

OpenStudy (anonymous):

its C

OpenStudy (holly00d1248):

Actually I think only one

OpenStudy (anonymous):

no i was answering the first one haha